OpenStudy (anonymous):

find the equation of the line through the given pair of points in standard form using only integers (5,-3) and (-2,4)

jimthompson5910 (jim_thompson5910):

what is the slope?

OpenStudy (anonymous):

y1-y2/x1-x2 ??

OpenStudy (anonymous):

-1

jimthompson5910 (jim_thompson5910):

slope: m = -1

jimthompson5910 (jim_thompson5910):

let's say we pick the point (5,-3) so x = 5 and y = -3

jimthompson5910 (jim_thompson5910):

y = mx+b -3 = -1*5 + b solve for b

OpenStudy (anonymous):

2

jimthompson5910 (jim_thompson5910):

the equation in slope intercept form is y = -1x + 2 or y = -x+2

jimthompson5910 (jim_thompson5910):

last step: turn that equation into standard form

OpenStudy (anonymous):

x+y+2

OpenStudy (anonymous):

=

OpenStudy (anonymous):

x+y=2

jimthompson5910 (jim_thompson5910):

x+y = 2 is correct

OpenStudy (anonymous):

but wait

jimthompson5910 (jim_thompson5910):

what's wrong?

OpenStudy (anonymous):

my multiple choices are different

OpenStudy (anonymous):

A) 7x+7y=14 B) 7x-7y=14 C)7x+7y=14 D)7x+7y=22

OpenStudy (anonymous):

are they wrong because the answer we came up with was the one i had origionaly

jimthompson5910 (jim_thompson5910):

ok we have x+y = 2 what can we do to both sides to get it to what they want

OpenStudy (anonymous):

times 7

jimthompson5910 (jim_thompson5910):

A and C are identical and you can simplify

jimthompson5910 (jim_thompson5910):

so we would have 7x+7y = 14 when we multiply both sides by 7

jimthompson5910 (jim_thompson5910):

which is either A or C

OpenStudy (anonymous):

c is -14

OpenStudy (anonymous):

the answer is a

OpenStudy (anonymous):

so thats what we have to do

jimthompson5910 (jim_thompson5910):

ok, then A it is

OpenStudy (anonymous):

ok thank you <3